annabanana:

Which is the graph of y – 3 = 2/3 (x + 6)?

Tranquility:

Hello! Welcome to QuestionCove! First distribute 2/3 using the distributive property a(b + c) = ab + ac a(b - c) = ab - ac Second, add 3 to both sides so that way you can isolate y You now have the equation in the form of y = mx + b Where m is the slope, and b is the y-intercept The y-intercept is the point (0, b) It is where the line intersects the y-axis (so the x-coordinate would be 0) We can graph it now that we know the equation. Since we already have one point, the y-intercept which is (0, b), first plot that. You can then use a test point by plugging in any value you wish for x, such as 1, and then solve for y. This is necessary because you need at least 2 points to draw a line. Let me know how far you get and/or if you need more help in understanding your question :)